OpenStudy (anonymous):

Please Help! Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

@satellite73 can you help?

OpenStudy (anonymous):

@Hero @phi ?

OpenStudy (phi):

do you know about "dot product" ?

OpenStudy (anonymous):

Yes but when i set it up i get a math error on my calculator

OpenStudy (anonymous):

unless i am multiplying them wrong or something?

OpenStudy (phi):

Forget the calculator for the moment. the dot product is defined two different ways: for \[ a= <a_x, a_y> \] and the same for b \[ a \cdot b = a_x b_x + a_y b_y \] also \[ a \cdot b = |a| |b| cos A\] where |a| is the "length" of a. same for b. A is the angle between the vectors a and b

OpenStudy (anonymous):

right so when i did that i got

OpenStudy (phi):

use the first definition to find \( a\cdot b\) use the 2nd definition to find \[ \cos A = \frac{a\cdot b}{|a| |b| } \]

OpenStudy (anonymous):

right so when i did that i got \[38\div \sqrt{20} \sqrt{73}\]

OpenStudy (anonymous):

I know its wrong but i dont get what i did incorrectly

OpenStudy (phi):

so you have \[ \cos A = \frac{38}{ \sqrt{20 \cdot 73} } \] (same as what you posted)

OpenStudy (anonymous):

Correct...

OpenStudy (anonymous):

Oh then i multiply them out! I thought i had to add the square roots. That gives me a final angle of 6 degrees right?

OpenStudy (phi):

yes. I get 6.0090 to the nearest tenth of a degree, that is 6.0 degrees

OpenStudy (anonymous):

Thank you so much!